OpenStudy (anonymous):

Use a calculator to find all solutions in the interval (0, 2π). Round the answers to two decimal places. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.) sec t = 9.44 Hint: The equation is equivalent to cos t = 1/(9.44).

OpenStudy (anonymous):

sec t = 9.44 1/cos t = 9.44 cos t = 1/9.44 t=arccos (1/9.44) Basic angle of t = 83.92 degrees. First quadrants and fourth quadrants , cos theta = +ve. t= 83.92, No solution for 2nd quadrants , No solution for 3rd quadrants , 360-83.92 t=83.92, 276.08